OpenStudy (anonymous):

need help finding perpendicular line of y=x+4 passing through (-7,1)

OpenStudy (anonymous):

y = m x + b y=-x+b using the given point: 1=-(-7)+b solve the above for b. then the perpendicular line is: y=-x-6 Refer to the attached plot.

OpenStudy (bradely):

Step by step solution http://www.mathskey.com/question2answer/4225/need-help